OpenStudy (anonymous):

A jacket usually sells for $57.00. If the jacket is 30% off, and sales tax is 7%, what is the total price of the jacket, including tax?

OpenStudy (mertsj):

P = .7(57)+.07(.7(57))

OpenStudy (anonymous):

\[57\times .7\times(1.07)\]

OpenStudy (anonymous):

30% of 57 is 17.1. 57 - 17.1 = 39.9. 7% of 39.9 is 2.79. 39.9 - 2.79 = 37.11. The jacket will cost a total of $37.11.

OpenStudy (mertsj):

Need to add the sales tax, Emmy, not subtract it.

OpenStudy (anonymous):

Whoops, the total will be $42.69. Thanks, Mertsj!